# The Status of Neutrino Cosmology: Standard $\Lambda$CDM, Extensions, and Tensions

### v2

@@ Abstract
We examine the performance of the six-parameter $\Lambda$CDM model and its extensions in light of recent cosmological observations, with particular focus on neutrino properties inferred from cosmology. Using a broad suite of nine combinations of datasets, with three separate analyses of the Planck Cosmic Microwave Background (CMB) data, and three separate supernovae (SNe) survey data, plus the recent DESI baryon acoustic oscillation (BAO) scale results, we derive constraints on the sum of neutrino masses ($\Sigma m_{\nu}$). Our results show upper limits in the range of $\Sigma m_{\nu}<76.9\,\mathrm{meV}$ to $\Sigma m_{\nu}<108\,\mathrm{meV}$ (95% CL). The variation in the limits on $\Sigma m_{\nu}$ arises from the separate analyses of the Planck CMB data and the separate supernovae datasets, as they relate to the inferred matter density and its relation to the sensitivity of the BAO scale and CMB lensing to $\Sigma m_{\nu}$. In the context of hierarchical mass models in $\Lambda$CDM, we find a $1.47\sigma$ preference for normal ordering (NO) over inverted ordering (IO), with similar values of preference across all datasets. Despite the strong constraints, an inclination towards massless neutrinos over NO remains weak at $1.36\sigma$. We find that a “negative” neutrino mass, inferred from the shape of the likelihood in the physical regime, $\Sigma m_{\nu}>0$, is only present at less than $2\sigma$. The origin of the strong $\Sigma m_{\nu}$ constraints arise primarily from the high CMB lensing signal, which disfavors the suppression of power from $\Sigma m_{\nu}$, and the DESI BAO scale, which are complementary to that from the suppression of power. We confirm that models allowing extra relativistic degrees of freedom, with $N_{\rm eff}\approx 3.5$, alleviate the Hubble tension. Significantly, we find a $3.3\sigma$ preference for a 0.1 eV partially thermalized sterile neutrino when the SH0ES $H_{0}$ measurement is included, a scale of interest in short-baseline oscillation experiment results. When $H_{0}$ is included, fully thermalized sterile neutrino models are as consistent as $\Lambda$CDM in fitting all datasets, but are disfavored otherwise. We also explore an 11-parameter model relaxing the dark energy equation of state and curvature, together with $\Sigma m_{\nu}$ and $N_{\rm eff}$, finding consistency with $\Lambda$CDM parameters except for the dark energy equation of state $w_{0}=-0.961^{+0.012}_{-0.037}$. Neutrino mass constraints in this extended model remain stringent, with $\Sigma m_{\nu}<97.0\,\mathrm{meV}$ (95% CL).
@@ I Introduction
where $\rho_{\nu}^{0}$ and $\rho_{\rm crit}^{0}$ are respectively the total neutrino and critical density of the Universe today and $h$ is the Hubble constant in units of $100\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$. In this expression, we use the temperature of neutrinos that provides $N_{\nu}=3.044$. The absolute neutrino mass scale has been probed using beta decay, with the KATRIN experiment currently providing the best bound on the beta-decay effective neutrino mass of $m_{\nu,\beta}<0.45\,\text{eV}$ at 90% confidence Aker et al. (2024). Despite these advancements, this bound is still about 20 times less stringent than state-of-the-art cosmology constraints Adame et al. (2024a).
@@ II.4 Type Ia SNe
It is important to note that the PP, U3 and DES Y5 datasets contain supernovae that overlap between both samples. To avoid any double counting, these datasets are never utilized together. For conciseness, we use the term “benchmark” to refer to the combination of the likelihoods from P18, CMB lensing, DESI DR1, and PP described above.
@@ IV.1 Hierarchical Neutrino Mass Ordering
$+\sqrt{m_{3}^{2}+|\Delta m_{\rm atm}^{2}|+\Delta m_{\rm solar}^{2}}\,.$
where the lightest neutrino mass eigenstate is $m_{1}$ for NO and $m_{3}$ for IO. A key question from cosmology is whether NO or IO is preferred by the data in the case of a hierarchical neutrino mass spectrum. That is, hierarchy is where $m_{1}\ll m_{2,3}$ (NO), or $m_{3}\ll m_{1,2}$. Assuming the minimal mass case for both orderings, i.e. $m_{1}=0$ for NO and $m_{3}=0$ for IO, we obtain the minimal $\Sigma m_{\nu}$ for NO and IO by using the observed mass-splittings. This defines our two massive neutrino cases as: NO with $\Sigma m_{\nu}=58\,\mathrm{meV}$ and IO with $\Sigma m_{\nu}=101\,\mathrm{meV}$.
@@
$\Delta\chi^{2}(\text{NO over IO})=2.16\left(1.47\sigma\right)\,,$ (6)
indicating a very weak preference for NO over IO. Considering the preference of the value inconsistent with neutrino physics of $\Sigma m_{\nu}=0$, and we find massless neutrinos preferred over NO by
$\Delta\chi^{2}(\Sigma m_{\nu}=0\,\text{over NO})=1.85\left(1.36\sigma\right)\,,$ (7)
again indicating very weak preference of $\Sigma m_{\nu}=0$ over NO. Importantly, the information on the matter content of the Universe from the supernovae surveys, which we adopt in our baseline cosmological dataset, slightly alleviates the cosmological constraint on neutrino mass. Without the PP SNe, we obtain $\Delta\chi^{2}(\text{NO over IO})=3.46\,(1.86\sigma)$ and $\Delta\chi^{2}(\Sigma m_{\nu}=0\,\text{over NO})=2.36\,(1.54\sigma)$. We find that the preference tests for the neutrino mass ordering are not substantially changed when adopting the CamSpec or Planck 2020 CMB data. To test the dependence on our default choice for modeling $\Sigma m_{\nu}$ as a single mass eigenstate, or as three degenerate states, we perform these $\chi^{2}$-tests with three degenerate neutrino masses, and find that the $\Delta\chi^{2}$ changes by less than 0.002 for both NO over IO and $\Sigma m_{\nu}=0$ over NO. Therefore, our results on evidence for mass ordering are not affected by our choice of modeling the hierarchy. Importantly, we find the sensitivity to neutrino mass ordering from P18+DESI+PP is about as strong as P18+Lensing+PP, showing that geometry is about as sensitive to neutrino mass as structure growth Loverde and Weiner (2024).
@@ IV.2 Free Neutrino Mass
There has been considerable interest recently on the tighter-than expected constraints on $\Sigma m_{\nu}$ to potentially indicate new physics, which could even be parameterized as “negative” neutrino mass by parametrizing the suppression of the growth of structure with massive neutrinos as enhanced growth when extending the growth of structure to be enhanced by the same linear suppression to linear enhancement of growth Craig et al. (2024); Green and Meyers (2024); Naredo-Tuero et al. (2024); Elbers et al. (2024); Allali and Notari (2024). To test the strength of the preference for “negative” neutrino mass, we fit Gaussian functions to the one-dimensional likelihoods of the chains for each dataset. We fit to all values of $[\Delta 2\ln({\mathcal{L}})]=2.5$ with respect to the likelihood maximum, as the MCMC process does not accurately sample the tails of the likelihood distributions. We recover the mean and standard deviation characteristic parameters of the associated Gaussian functions for the different datasets and calculate the deviation in $\sigma$ from the central value of the Gaussian to the neutrino physics minimal-mass NO case, i.e. $\Sigma m_{\nu}=58\,\mathrm{meV}$. For all of the datasets we have analyzed where $\Sigma m_{\nu}$ is free, we find the deviation of the NO from the mean of the inferred likelihood to be between $1.13\sigma$ and $1.97\sigma$, always remaining below $2\sigma$.
@@ V.2.2 Fully Thermalized
V.3 The $S_{8}$ Tension and Massive Active or Sterile Neutrinos
[…]
Parameters Benchmark Benchmark + $H_{0}$ Post. + $S_{8}$ Post. + $H_{0}$ $S_{8}$ Post.

### v3

# Status of neutrino cosmology: Standard $\Lambda$CDM, extensions, and tensions

@@ Abstract
We examine the performance of the six-parameter $\Lambda$CDM model and its extensions in light of recent cosmological observations, with particular focus on neutrino properties inferred from cosmology. Using a broad suite of nine combinations of datasets, with three separate analyses of the Planck Cosmic Microwave Background (CMB) data, and three separate supernovae (SNe) survey data, plus the recent DESI baryon acoustic oscillation (BAO) scale results, we derive constraints on the sum of neutrino masses ($\Sigma m_{\nu}$). Our results show upper limits in the range of $\Sigma m_{\nu}<76.9\,\mathrm{meV}$ to $\Sigma m_{\nu}<108\,\mathrm{meV}$ (95% CL). The variation in the limits on $\Sigma m_{\nu}$ arises from the separate analyses of the Planck CMB data and the separate supernovae datasets, as they relate to the inferred matter density and its relation to the sensitivity of the BAO scale and CMB lensing to $\Sigma m_{\nu}$. In the context of hierarchical mass models in $\Lambda$CDM, we find a $1.47\sigma$ preference for normal ordering (NO) over inverted ordering (IO), with similar values of preference across all datasets. Despite the strong constraints, an inclination towards the non-standard massless neutrinos over NO remains weak at $1.36\sigma$. We find that a “negative” neutrino mass, inferred from the shape of the likelihood in the physical regime, $\Sigma m_{\nu}>0$, is only present at less than $2\sigma$. The origin of the strong $\Sigma m_{\nu}$ constraints arise primarily from the high CMB lensing signal, which disfavors the suppression of power from $\Sigma m_{\nu}$, and the DESI BAO scale, which are complementary to that from the suppression of power. We confirm that models allowing extra relativistic degrees of freedom, with $N_{\rm eff}\approx 3.5$, alleviate the Hubble tension. Significantly, we find a $3.3\sigma$ preference for a 0.1 eV partially thermalized sterile neutrino when the SH0ES $H_{0}$ measurement is included, a scale of interest in short-baseline oscillation experiment results. When $H_{0}$ is included, fully thermalized sterile neutrino models are as consistent as $\Lambda$CDM in fitting all datasets, but are disfavored otherwise. We also explore an 11-parameter model relaxing the dark energy equation of state and curvature, together with $\Sigma m_{\nu}$ and $N_{\rm eff}$, finding consistency with $\Lambda$CDM parameters except for the dark energy equation of state $w_{0}=-0.961^{+0.012}_{-0.037}$. Neutrino mass constraints in this extended model remain stringent, with $\Sigma m_{\nu}<97.0\,\mathrm{meV}$ (95% CL).
@@ I Introduction
where $\rho_{\nu}^{0}$ and $\rho_{\rm crit}^{0}$ are respectively the total neutrino and critical density of the Universe today and $h$ is the Hubble constant in units of $100\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$. In this expression, we use the temperature of neutrinos that provides $N_{\nu}=3.044$. The absolute neutrino mass scale has been probed using beta decay, with the KATRIN experiment currently providing the best bound on the beta-decay effective neutrino mass of $m_{\nu,\beta}<0.45\,\text{eV}$ at 90% confidence Aker et al. (2024).
@@ II.4 Type Ia SNe
It is important to note that the PP, U3 and DES Y5 datasets contain supernovae that overlap between the samples. To avoid any double counting, these datasets are never utilized together. Table 1 summarizes all the datasets that we have just defined. For conciseness, we use the term “benchmark” to refer to the combination of the likelihoods from P18, CMB lensing, DESI DR1, and PP described above.
@@ IV.1 Hierarchical Neutrino Mass Ordering
$+\sqrt{m_{3}^{2}+|\Delta m_{\rm atm}^{2}|+\Delta m_{\rm solar}^{2}}\,,$
where the lightest neutrino mass eigenstate is $m_{1}$ for NO and $m_{3}$ for IO. A key question from cosmology is whether NO or IO is preferred by the data in the case of a hierarchical neutrino mass spectrum. That is, hierarchy is where $m_{1}\ll m_{2,3}$ (NO), or $m_{3}\ll m_{1,2}$ (IO). Assuming the minimal mass case for both orderings, i.e. $m_{1}=0$ for NO and $m_{3}=0$ for IO, we obtain the minimal $\Sigma m_{\nu}$ for NO and IO by using the observed mass-splittings. This defines our two massive neutrino cases as: NO with $\Sigma m_{\nu}=58\,\mathrm{meV}$ and IO with $\Sigma m_{\nu}=101\,\mathrm{meV}$.
@@
$\Delta\chi^{2}(\text{NO over IO})=2.16\,,$ (6)
indicating a very weak preference, approximately $1.47\sigma$, for NO over IO. Considering the preference of the value inconsistent with neutrino physics of $\Sigma m_{\nu}=0$, and we find non-standard massless neutrinos preferred over NO by
$\Delta\chi^{2}(\Sigma m_{\nu}=0\,\text{over NO})=1.85\,,$ (7)
indicating very weak preference, approximately $1.36\sigma$, of $\Sigma m_{\nu}=0$ over NO. Importantly, the information on the matter content of the Universe from the supernovae surveys, which we adopt in our baseline cosmological dataset, slightly alleviates the cosmological constraint on neutrino mass. Without the PP SNe, we obtain $\Delta\chi^{2}(\text{NO over IO})=3.46\,(1.86\sigma)$ and $\Delta\chi^{2}(\Sigma m_{\nu}=0\,\text{over NO})=2.36\,(1.54\sigma)$. We find that the preference tests for the neutrino mass ordering are not substantially changed when adopting the CamSpec or Planck 2020 CMB data. To test the dependence on our default choice for modeling $\Sigma m_{\nu}$ as a single mass eigenstate, or as three degenerate states, we perform these $\chi^{2}$-tests with three degenerate neutrino masses, and find that the $\Delta\chi^{2}$ changes by less than 0.002 for both NO over IO and $\Sigma m_{\nu}=0$ over NO. Therefore, our results on evidence for mass ordering are not affected by our choice of modeling the hierarchy. Importantly, we find the sensitivity to neutrino mass ordering from P18+DESI+PP is about as strong as P18+Lensing+PP, showing that geometry is about as sensitive to neutrino mass as structure growth Loverde and Weiner (2024).
@@ IV.2 Free Neutrino Mass
There has been considerable interest recently on the tighter-than expected constraints on $\Sigma m_{\nu}$ to potentially indicate new physics, which could even be parameterized as “negative” neutrino mass by parameterizing the suppression of the growth of structure with massive neutrinos as enhanced growth when extending the growth of structure to be enhanced by the same linear suppression to linear enhancement of growth Craig et al. (2024); Green and Meyers (2024); Naredo-Tuero et al. (2024); Elbers et al. (2024); Allali and Notari (2024). To test the strength of the preference for “negative” neutrino mass, we fit Gaussian functions to the one-dimensional likelihoods of the chains for each dataset. We fit to all values of $[\Delta 2\ln({\mathcal{L}})]=2.5$ with respect to the likelihood maximum, as the MCMC process does not accurately sample the tails of the likelihood distributions. We recover the mean and standard deviation characteristic parameters of the associated Gaussian functions for the different datasets and calculate the deviation in $\sigma$ from the central value of the Gaussian to the neutrino physics minimal-mass NO case, i.e. $\Sigma m_{\nu}=58\,\mathrm{meV}$. For all of the datasets we have analyzed where $\Sigma m_{\nu}$ is free, we find the deviation of the NO from the mean of the inferred likelihood to be between $1.13\sigma$ and $1.97\sigma$, always remaining below $2\sigma$.
@@ V.2.2 Fully Thermalized
Parameters Benchmark Benchmark + $H_{0}$ Posterior + $S_{8}$ Posterior + $H_{0}$ $S_{8}$ Posteriors